# Energy, enstrophy and helicity transfers in polymeric turbulence

### v1

@@ Abstract
We characterise the scale-by-scale transfers of energy, enstrophy and helicity in homogeneous and isotropic polymeric turbulence using direct numerical simulations. The microscale Reynolds number is set to $Re_{\lambda}\approx 460$, and the Deborah number $De=\tau_{p}/\tau_{f}$ is varied between $1/9\leq De\leq 9$; $\tau_{p}$ is the polymeric relaxation time and $\tau_{f}$ is the turnover time of the largest scales of the flow. The study relies on the exact scale-by-scale budget equations (derived from the the governing model equations) for energy, enstrophy and helicity, which account for the back-reaction of the polymers on the flow. Polymers act as a sink/source in the flow, and provide alternative routes for the scale-by-scale transfers of the three quantities, whose relevance changes with $De$. We find that polymers deplete the nonlinear energy cascade mainly at smaller scales, by weakening both the extreme forward as well as reverse local events. The new polymer-driven energy flux dominates at small scales for $De\geq 1$, and on average transfers energy from larger to smaller scales with localised backscatter events. Polymers weaken the stretching of vorticity with the enstrophy being mainly generated by the fluid-polymer interaction, especially when $De\geq 1$. Accordingly, an inspection of the small-scale flow topology shows that polymers favour events with two-dimensional state of straining, and promote/inhibit extreme extension/rotation events: in polymeric turbulence shear and planar extensional flows are more probable. The helicity injected at the largest scales shows a similar transfer process to as energy, being mainly driven by the nonlinear cascade at large scales and by the polymer-driven flux at small scales. Polymers are found to favour events that break the small-scale mirror symmetry, with the relative helicity monotonically increasing with $De$ at all scales.
@@ 1 Introduction
A small concentration of polymers in a turbulent flow results in a substantial decrease of drag, as is well documented in several experimental, numerical and theoretical works (Toms, 1948; Lumley, 1969; Berman, 1978; Benzi & Ching, 2018). In a turbulent wall-bounded flow, the polymeric chains influence the energy- momentum transfers in the real space and in the space of scales, giving rise to an increased buffer region that eventually leads to drag reduction (see for example Warwaruk & Ghaemi, 2024). It is known that large drag reduction may occur without a substantial reduction of the turbulent kinetic energy, which may indeed be either smaller or larger compared to the purely Newtonian counterpart (Tsinober, 1990). Polymeric turbulence, therefore, is not necessarily associated with the suppression of turbulence, but rather with qualitative/quantitative changes of its structure. Despite the large interest, however, the way polymeric additives modulate turbulence is not completely understood yet, and their influence on the statistical nature of the velocity fluctuations is still under debate. In this work, we consider the idealised set up of homogeneous and isotropic turbulence (HIT), and investigate the influence of polymeric additives on the cascades of energy, helicity and enstrophy in the space of scales.
@@ 1.1 Polymers in homogeneous and isotropic turbulence
Two dimensionless numbers are needed to characterise turbulent flows of dilute polymeric solutions in a triperiodic box. They are the Reynolds number $Re$, which estimates the importance of the inertial term compared to the viscous one in the Navier–Stokes equations, and the Deborah number $De$ that quantifies the unique polymer relaxation time-scale with respect to the large time scale of the flow. A large De therefore means the polymers are more elastic, and extensible. Turbulent drag reduction has been mainly observed for large Re and De, with an evident departure from the Kolmogorov predictions already in the idealised setting of HIT (Tabor & Gennes, 1986; Bhattacharjee & Thirumalai, 1991; Fouxon & Lebedev, 2003). At rather low Re of $Re_{\lambda}=u^{\prime}\lambda/\nu\approx 80$ ($u^{\prime}$ is the typical velocity fluctuation, $\lambda$ is the Taylor length-scale and $\nu$ is the kinematic fluid viscosity), Perlekar et al. (2010) found that in the presence of the polymers the energy content decreases at the intermediate scales and significantly increases at the smallest scales, in the deep dissipative range. They also observed that the fluid dissipation monotonically decreases as $De$ increases, although they fixed the injected energy using the forcing introduced by Lamorgese et al. (2005), hinting that polymers give origin to an alternative route to the classical energy cascade.
@@
At large $Re$ and $De$, the polymer additives modify the nature of the energy cascade and result in a significant alteration of the energy distribution amongst scales. For large enough separation between the energy injection scale $\mathcal{L}$ and the Kolmogorov scale $\eta$, there exists an intermediate scale $r_{p}$ that separates two different regimes. For $r_{p}\leq r\leq\mathcal{L}$ the energy cascade resembles that of a Newtonian flow, with the classical second-order structure function scaling $S_{2}(r)\sim r^{2/3}$ predicted by the Kolmogorov theory. For $\eta\leq r\leq r_{p}$, the elastic range of scales, energy is transferred among scales by the polymeric microstructure as well, and the second-order structure function $S_{2}$ increases faster than $r^{2/3}$, deviating from the Kolmogorov predictions (De Angelis et al., 2005). Recently, Zhang et al. (2021) and Rosti et al. (2023) provided experimental and numerical evidence that, for large enough $De$, there is a range of scales $\eta\lesssim r\lesssim r_{p}$ where $S_{2}(r)\sim r^{\xi}$, with $\xi\approx 1.3\approx 4/3$. By studying the scale-by-scale energy budget in Fourier space, Rosti et al. (2023) observed that the energy flux at only the large scales is dominated by the advective nonlinearity for large enough $De$ (as for purely Newtonian turbulence), and that the polymeric stresses become dominant at smaller scales. Interestingly, they also observed that $r_{p}$ has a non-monotonic dependence on $De$, and follows the cross-over scale between the nonlinear flux and the non-Newtonian flux. As a consequence, the width of the elastic range of scales does not show a monotonic dependence on the $De$, and is maximum for $De\approx 1$. More recently, Singh & Rosti (2024) investigated how polymeric turbulence depends on $De$ and $Re$, linking the low-$Re$ regime described by Singh et al. (2024) with the large-$Re$ regime described by Rosti et al. (2023). They found that in polymeric turbulence the dissipation field is intermittent and shows a qualitative similar distribution as in purely Newtonian turbulence.
@@ 1.2 Enstrophy and Helicity
In HIT, enstrophy is well known to be closely related to dissipation: they have the same mean value up to a constant given by the fluid viscosity $\nu$ (Tsinober, 2001). However, they are two different descriptors of the structure of the small scales of turbulence (Sreenivasan & Antonia, 1997), and extreme events of dissipation and enstrophy are rather very different. Extreme dissipation represents intense local strain, while large enstrophy represents strong vortical motions. Despite their intrinsic difference, however, Yeung et al. (2012) showed that at sufficiently large $Re$ extreme events of dissipation and enstrophy scale similar, and tend to occur together. Enstrophy shares some common features with energy as well, such as a transfer from large to small scales, with the dissipation being confined at the small scales. However, the conventional notion of a cascade does not apply for enstrophy: it is not inviscidly conserved because vortex stretching ($\bm{\omega}\cdot\bm{\nabla}\bm{u}$) acts as a source of vorticity at all scales (Davidson et al., 2008). Most of enstrophy is generated at the small scales and does not have to be transferred among space and scales to be dissipated. Nevertheless, vortex stretching is active also in the inertial range of scales (where dissipation is negligible), and this enstrophy has to be transferred from large to small scales to be eventually destroyed. The importance of vortex stretching in turbulence is well known since the pioneering works of Taylor (1938); Betchov (1956); Ashurst et al. (1987). Though, there is a lack of consensus of the role of vortex stretching in the energy cascade. Despite the earlier works (see for example Davidson, 2004; Davidson et al., 2008; Doan et al., 2018) providing evidence about the relation between energy cascade and vortex stretching, some recent works have shown that the energy cascade is mainly driven by the self-amplification of the strain-rate field, and that vortex stretching plays a key role only during fluctuations of the cascade about its average value (Carbone & Bragg, 2020).
@@
Helicity, a pseudoscalar defined as $h=\bm{\omega}\cdot\bm{u}$, is another inviscid invariant of the three-dimensional ($3D$) Navier–Stokes equations. That is, like energy, helicity is also conserved by the non linearity of the Navier–Stokes equations (Moffatt, 1969). It admits topological interpretations in relation to the linkages of the flow vortex lines (Moffatt, 1969; Moffatt & Tsinober, 1992): it is related to the knottedness of the vorticity lines, and is a descriptor of the breaking of parity invariance (mirror symmetry). Unlike energy, however, it is not a sign-definite quantity. The invariance of helicity has been associated with the conservation of the linkages of the vortex lines that move with the flow (Moffatt, 1969), and it is based on two main notions, i.e. (i) vortex lines behave like material lines under evolution of the inviscid Navier–Stokes equations, and (ii) the flux of vorticity through any open surface bounded by a curve moving with the fluid is conserved. Based on the conservation of total helicity, Brissaud et al. (1973) were the first to envisage the possible simultaneous existence of energy and helicity cascades in $3D$ turbulence, similar to what happens in two-dimensional ($2D$) turbulence where the two invariants are the energy and the enstrophy (Boffetta & Ecke, 2012; Falkovich et al., 2017). Based on phenomenological arguments, they proposed that two different scenarios are possible, when dealing with helical flows. The first admits a simultaneous cascade of energy and helicity that leads to a $-5/3$ power-law spectrum for both quantities. The second scenario of a pure helicity cascade (i.e. no energy cascade) predicts a range of power-laws for both energy and helicity. Kraichnan (1973) argued that, unlike in $2D$ turbulence where the existence of invariant enstrophy effectively blocks the forward energy cascade, the possibility of a joint energy and helicity cascade is more plausible (see also André & Lesieur, 1977) in $3D$ turbulence. This was later confirmed by Polifke & Shtilman (1989), and by the numerical simulations of Borue & Orszag (1997). They observed a cascade of helicity from large to small scales, and did not detect an inverse cascade of helicity. In the following years, several authors studied helicity cascade in helical homogeneous isotropic turbulence. Chen et al. (2003b) found using direct numerical simulations that the energy and helicity fluxes feature a plateau in the inertial range of scales confirming the existence of a joint helicity and energy cascade from large to small scales. They also investigated intermittency in these fluxes and observed that the scaling exponents for the helicity fluxes are smaller compared to those for energy. The helicity flux is thus intrinsically more intermittent than the energy flux, consistent with the observation that helicity essentially behaves like a passive scalar (Sreenivasan & Antonia, 1997; Romano & Antonia, 2001). Over the last years, several authors have studied helicity cascades, mainly using the Fourier space statistics (see for example Biferale et al., 2013; Alexakis, 2017; Pouquet et al., 2019). Most of these works exploit an exact decomposition of the velocity field in a helical Fourier basis to properly account for triad interaction between wavenumbers (Constantin & Majda, 1988; Waleffe, 1992).
@@ 1.3 The Kármán-Howarth-Hill or the Generalised Kolmogorov Equation
It is therefore clear that to fully characterise the influence of the polymeric additives on the Kolmogorov picture of turbulence, one has to consider also enstrophy and helicity besides the turbulent kinetic energy. In this respect, Baj et al. (2022) introduced a generalised framework which is suitable for investigating the scale-space transfers of energy, enstrophy and helicity in Newtonian turbulent flows. Their framework extends the generalised Kolmogorov equation, or GKE, (Marati et al., 2004; Danaila et al., 2004; Cimarelli et al., 2013; Gatti et al., 2020; Gattere et al., 2023), also known as Kármán-Howarth-Monin-Hill or Kármán-Howarth-Hill equation (Alves Portela et al., 2017; Yasuda & Vassilicos, 2018; Alves Portela et al., 2020; Yao et al., 2024), introduced by Hill (2001, 2002). The GKE is a generalisation of the Kármán-Howarth equation (Frisch, 1995; Pope, 2000), and is directly derived from the Navier–Stokes equations without any assumptions; it does not require either isotropy nor homogeneity. The GKE is an exact budget equation for the second-order structure function, commonly referred to as scale energy (Davidson & Pearson, 2005), and characterises the mechanisms of production, transfer and dissipation of energy in the combined space of scales and positions. The GKE and its generalisations have been applied to several flow configurations over the years, mainly to study how inhomogeneity changes the Richardson and Kolmogorov picture of turbulence (Casciola et al., 2003; Cimarelli et al., 2016; Alves Portela et al., 2017; Mollicone et al., 2018; Cimarelli et al., 2021; Chiarini et al., 2022b, a; Apostolidis et al., 2023). De Angelis et al. (2005) used the GKE to investigate the influence of polymeric additives on the energy cascade in homogeneous isotropic turbulence, but at rather small Deborah numbers ($De\leq 0.5$), for which the influence of the polymeric stresses is subdominant at all scales.
@@ 1.4 The present study
In this work we investigate the influence of the polymeric additives on the scale transfers of energy, helicity and enstrophy in HIT at a relatively large Re and over a wide range of De. First, we extend the formulation of Baj et al. (2022) and provide the exact scale-by-scale budget equations for energy, helicity and enstrophy for polymer-laden turbulent flows. Compared to the purely Newtonian case, the new resulting budget equations feature additional terms that capture the influence of the polymers on the scale-space transfers, and detail the scale-space exchanges of energy, helicity and enstrophy between the fluid and polymeric phases. The new set of equations has been derived without any approximation, and are valid for a generic inhomogeneous and anisotropic turbulent flow. Then, we use these equations to elucidate the influence of the polymers on the production and transfers of energy, helicity and enstrophy in a homogeneous isotropic turbulent flow of dilute polymeric suspensions. New insights are provided, with a particular look at the influence of the polymers on the intermittent nature of the transfers, and on the local flow topology. The study relies on the database introduced by Singh & Rosti (2024) and obtained by means of direct numerical simulations. The Reynolds number is set at $Re_{\lambda}\approx 460$, while the Deborah number is varied in the $1/9\leq De\leq 9$ range.
The remainder of the work is structured as follows. In §2 we introduce the scale-space budget equations. In §2.3 the details of the DNS database used in this work are briefly recalled. Sections §4, §5 and §6 deal with our findings, and are respectively devoted to the scale-by-scale budget equations for energy, enstrophy and helicity. In §5 the effect of the polymers on the local flow topology is also discussed in relation with their influence on the vortex stretching. A concluding discussion and perspectives are then provided in §7.

### v2

@@ Abstract
We investigate the scale-by-scale transfers of energy, enstrophy and helicity in homogeneous and isotropic polymeric turbulence using direct numerical simulations. The study relies on the exact scale-by-scale budget equations, derived from the governing model equations, that fully capture the back-reaction of polymers on the fluid dynamics. Polymers act as dynamic sinks and sources and open alternative routes for interscale transfer whose significance is modulated by their elasticity, quantified through the Deborah number ($De$). Polymers primarily deplete the nonlinear energy cascade at small scales, by attenuating intense forward and inverse transfer events. At sufficiently high $De$, a polymer-driven flux emerges and dominates at small scales, transferring on average energy from larger to smaller scales, while allowing for localised backscatter. For enstrophy, polymers inhibit the stretching of vorticity, with fluid-polymer interactions becoming the primary enstrophy source at high $De$. Accordingly, an analysis of the small-scale flow topology reveals that polymers promote two-dimensional straining states and enhance the occurrence of shear and planar extensional flows, while suppressing extreme rotation events. Helicity, injected at large scales, exhibits a transfer mechanism analogous to energy, being dominated by nonlinear dynamics at large scales and by polymer-induced fluxes at small scales. Polymers enhance the breakdown of small-scale mirror symmetry, as indicated by a monotonic increase in relative helicity with $De$ across all scales.
@@ 1 Introduction
A small concentration of polymers in a turbulent flow results in a substantial decrease of drag, as well documented in several experimental, numerical and theoretical works (Toms, 1948; Lumley, 1969; Berman, 1978; Benzi & Ching, 2018). In a turbulent wall-bounded flow, the polymeric chains influence the energy- momentum transfers in the real space and in the space of scales, giving rise to an increased buffer region that eventually leads to drag reduction (see for example Warwaruk & Ghaemi, 2024). It is known that large drag reduction may occur without a substantial reduction of the turbulent kinetic energy, which may indeed be either smaller or larger compared to the purely Newtonian counterpart (Tsinober, 1990). Polymeric turbulence, therefore, is not necessarily associated with the suppression of turbulence, but rather with qualitative/quantitative changes of its structure. Despite the large interest, however, the way polymeric additives modulate turbulence is not completely understood yet, and their influence on the statistical nature of the velocity fluctuations is still under debate. In this work, we consider the idealised set up of homogeneous and isotropic turbulence (HIT), and investigate the influence of polymeric additives on the cascades of energy, helicity and enstrophy in the space of scales.
@@ 1.1 Polymers in homogeneous and isotropic turbulence
Two dimensionless numbers are needed to characterise turbulent flows of dilute polymeric solutions in a triperiodic box. They are the Reynolds number $Re$, which estimates the importance of the inertial term compared to the viscous one in the Navier–Stokes equations, and the Deborah number $De$ that quantifies the unique polymer relaxation time scale with respect to the large time scale of the flow. A large $De$ therefore means the polymers are more elastic, and extensible. Turbulent drag reduction has been mainly observed for large Re and $De$, with an evident departure from the Kolmogorov predictions already in the idealised setting of HIT (Tabor & Gennes, 1986; Bhattacharjee & Thirumalai, 1991; Fouxon & Lebedev, 2003). At a relatively low Reynolds number, $Re_{\lambda}=u^{\prime}\lambda/\nu\approx 80$ ($u^{\prime}$ is the typical velocity fluctuation, $\lambda$ is the Taylor length-scale and $\nu$ is the kinematic fluid viscosity), Perlekar et al. (2010) observed that in the presence of the polymers the energy content decreases at the intermediate scales and significantly increases at the smallest scales, in the deep dissipative range. They also observed that the fluid dissipation monotonically decreases as $De$ increases, although they fixed the injected energy using the forcing introduced by Lamorgese et al. (2005), hinting that polymers give origin to an alternative route to the classical energy cascade.
@@
At large $Re$ and $De$, the polymer additives modify the nature of the energy cascade and result in a significant alteration of the energy distribution amongst scales. For large enough separation between the energy injection scale $\mathcal{L}$ and the Kolmogorov scale $\eta$, there exists an intermediate scale $r_{p}$ that separates two different regimes. For $r_{p}\leq r\leq\mathcal{L}$ the energy cascade resembles that of a Newtonian flow, with the classical second-order structure function scaling $S_{2}(r)\sim r^{2/3}$ predicted by the Kolmogorov theory. For $\eta\leq r\leq r_{p}$, the elastic range of scales, energy is transferred among scales by the polymeric microstructure as well, and the second-order structure function $S_{2}$ increases faster than $r^{2/3}$, deviating from the Kolmogorov predictions (De Angelis et al., 2005). Recently, Zhang et al. (2021) and Rosti et al. (2023) provided experimental and numerical evidence that, for large enough $De$, there is a range of scales $\eta\lesssim r\lesssim r_{p}$ where $S_{2}(r)\sim r^{\xi}$, with $\xi\approx 1.3\approx 4/3$. By analysing the scale-by-scale energy budget in Fourier space, Rosti et al. (2023) showed that, for sufficiently large $De$, the energy flux remains dominated by advective nonlinearities at the large scales as in Newtonian turbulence, while polymeric stresses become the dominant contribution at smaller scales. Interestingly, they also observed that $r_{p}$ has a non-monotonic dependence on $De$, and follows the cross-over scale between the nonlinear flux and the polymer-driven flux. As a consequence, the width of the elastic range of scales does not show a monotonic dependence on the $De$, and is maximum for $De=\mathcal{O}(1)$. More recently, Singh & Rosti (2025) investigated how polymeric turbulence depends on $De$ and $Re$, linking the low-$Re$ regime described by Singh et al. (2024) with the large-$Re$ regime described by Rosti et al. (2023). They found that in polymeric turbulence the dissipation field is intermittent and shows a qualitative similar distribution as in purely Newtonian turbulence.
@@ 1.2 Enstrophy and Helicity
In HIT, enstrophy is well known to be closely related to dissipation: they have the same mean value up to a constant given by the fluid viscosity $\nu$ (Tsinober, 2001). However, they are two different descriptors of the structure of the small scales of turbulence (Sreenivasan & Antonia, 1997), and extreme events of dissipation and enstrophy are rather different. Extreme dissipation represents intense local strain, while large enstrophy represents strong vortical motions. Despite their intrinsic difference, however, Yeung et al. (2012) showed that at sufficiently large $Re$ extreme events of dissipation and enstrophy scale similar, and tend to occur together. Enstrophy shares some common features with energy as well, such as a transfer from large to small scales, with the dissipation being confined at the small scales. However, the conventional notion of a cascade does not apply for enstrophy: it is not inviscidly conserved because vortex stretching ($\bm{\omega}\cdot\bm{\nabla}\bm{u}$) acts as a source of vorticity at all scales (Davidson et al., 2008). Most of enstrophy is generated at the small scales and does not have to be transferred among space and scales to be dissipated. Nevertheless, vortex stretching is active also in the inertial range of scales (where dissipation is negligible), and this enstrophy has to be transferred from large to small scales to be eventually destroyed. The importance of vortex stretching in turbulence is well known since the pioneering works of Taylor (1938); Betchov (1956); Ashurst et al. (1987). Though, there is a lack of consensus of the role of vortex stretching in the energy cascade. Despite the earlier works (see for example Davidson, 2004; Davidson et al., 2008; Doan et al., 2018) providing evidence about the relation between energy cascade and vortex stretching, some recent works have shown that the energy cascade is mainly driven by the self-amplification of the strain-rate field, and that vortex stretching plays a key role only during fluctuations of the cascade about its average value (Carbone & Bragg, 2020; Johnson, 2020).
@@
Helicity, a pseudoscalar defined as $h=\bm{\omega}\cdot\bm{u}$, is another inviscid invariant of the three-dimensional ($3D$) Navier–Stokes equations. That is, like energy, helicity is also conserved by the non linearity of the Navier–Stokes equations (Moffatt, 1969). It admits topological interpretations in relation to the linkages of the flow vortex lines (Moffatt, 1969; Moffatt & Tsinober, 1992): it is related to the knottedness of the vorticity lines, and is a descriptor of the breaking of parity invariance (mirror symmetry). Unlike energy, however, it is not a sign-definite quantity. The invariance of helicity has been associated with the conservation of the linkages of the vortex lines that move with the flow (Moffatt, 1969), and it is based on two main notions, i.e. (i) vortex lines behave like material lines under evolution of the inviscid Navier–Stokes equations, and (ii) the flux of vorticity through any open surface bounded by a curve moving with the fluid is conserved. Based on the conservation of total helicity, Brissaud et al. (1973) were the first to envisage the possible simultaneous existence of energy and helicity cascades in $3D$ turbulence, similar to what happens in two-dimensional ($2D$) turbulence where the two invariants are the energy and the enstrophy (Boffetta & Ecke, 2012; Falkovich et al., 2017). Based on phenomenological arguments, they proposed that two different scenarios are possible, when dealing with helical flows. The first admits a simultaneous cascade of energy and helicity that leads to a 5/3-5/3 power-law spectrum for both quantities. The second scenario of a pure helicity cascade (i.e. no energy cascade) predicts a range of power-laws for both energy and helicity. Kraichnan (1973) argued that, unlike in $2D$ turbulence where the existence of invariant enstrophy effectively blocks the forward energy cascade, the possibility of a joint energy and helicity cascade is more plausible (see also André & Lesieur, 1977) in $3D$ turbulence. This was later confirmed by Polifke & Shtilman (1989), and by the numerical simulations of Borue & Orszag (1997). They observed a cascade of helicity from large to small scales, and did not detect any inverse cascade. In the following years, several authors studied helicity cascade in helical homogeneous isotropic turbulence. Chen et al. (2003b) found, using direct numerical simulations, that the energy and helicity fluxes feature a plateau in the inertial range of scales confirming the existence of a joint helicity and energy cascade from large to small scales. They also investigated intermittency in these fluxes and observed that the scaling exponents for the helicity fluxes are smaller compared to those for energy. The helicity flux is thus intrinsically more intermittent than the energy flux, consistent with the observation that helicity essentially behaves like a passive scalar (Sreenivasan & Antonia, 1997; Romano & Antonia, 2001). Over the last years, several authors have studied helicity cascades, mainly using the Fourier space statistics (see for example Biferale et al., 2013; Alexakis, 2017; Pouquet et al., 2019). Most of these works exploit an exact decomposition of the velocity field in a helical Fourier basis to properly account for triad interaction between wavenumbers (Constantin & Majda, 1988; Waleffe, 1992).
@@ 1.3 The Kármán-Howarth-Hill or the Generalised Kolmogorov Equation
It is therefore clear that to fully characterise the influence of the polymeric additives on the Kolmogorov picture of turbulence, one has to consider also enstrophy and helicity besides the turbulent kinetic energy. In this respect, Baj et al. (2022) introduced a generalised framework which is suitable for investigating the scale-space transfers of energy, enstrophy and helicity in Newtonian turbulent flows. Their framework extends the generalised Kolmogorov equation, or GKE, (Marati et al., 2004; Danaila et al., 2004; Cimarelli et al., 2013; Gatti et al., 2020; Gattere et al., 2023), also known as Kármán-Howarth-Monin-Hill or Kármán-Howarth-Hill equation (Alves Portela et al., 2017; Yasuda & Vassilicos, 2018; Alves Portela et al., 2020; Yao et al., 2024b), introduced by Hill (2001, 2002). The GKE is a generalisation of the Kármán-Howarth equation (Frisch, 1995; Pope, 2000), and is directly derived from the Navier–Stokes equations without any assumptions; it does not require either isotropy nor homogeneity. The GKE is an exact budget equation for the second-order structure function, commonly referred to as scale energy (Davidson & Pearson, 2005), and characterises the mechanisms of production, transfer and dissipation of energy in the combined space of scales and positions. The GKE and its generalisations have been applied to several flow configurations over the years, mainly to study how inhomogeneity changes the Richardson and Kolmogorov picture of turbulence (Casciola et al., 2003; Cimarelli et al., 2016; Alves Portela et al., 2017; Mollicone et al., 2018; Cimarelli et al., 2021; Chiarini et al., 2022b, a; Apostolidis et al., 2023). De Angelis et al. (2005) used the GKE to investigate the influence of polymeric additives on the energy cascade in homogeneous isotropic turbulence, but at rather small Deborah numbers ($De\leq 0.5$), for which the influence of the polymeric stresses is subdominant at all scales.
@@ 1.4 The present study
In this work we investigate the influence of the polymeric additives on the scale transfers of energy, helicity and enstrophy in homogeneous and isotropic turbulence at a relatively large Re and over a wide range of $De$. First, we extend the formulation of Baj et al. (2022) and provide the exact budget equations for the scale energy, helicity and enstrophy for polymer-laden turbulent flows. Compared to the purely Newtonian case, the resulting budget equations incorporate additional terms that represent sources and sinks stemming from fluid–polymer interactions and capture the polymers’ influence on transfers both across scales and along the flow’s inhomogeneous directions. The new set of equations has been derived without any approximation, and are valid for a generic inhomogeneous and anisotropic turbulent flow. Then, we use these equations to elucidate the influence of the polymers on the production, transfer and dissipation of energy, helicity and enstrophy in a homogeneous isotropic turbulent flow of dilute polymeric suspensions. New insights are provided, with a particular look at the influence of the polymers on the intermittent nature of the transfers, and on the local flow topology. The study relies on the database introduced by Singh & Rosti (2025) and obtained by means of direct numerical simulations. The Reynolds number is set at $Re_{\lambda}\approx 460$, while the Deborah number is varied in the $1/9\leq De\leq 9$ range.
The remainder of the work is structured as follows. In §2 we introduce the budget equations. Sections §3, §4 and §5 deal with our findings, with §4.1, §4.2 and §4.3 being respectively devoted to energy, enstrophy and helicity scale budgets. §5 examines the influence of polymers on the local flow topology, particularly in relation to their effect on the scale-space budgets. A concluding discussion and perspectives are eventually provided in §6.